OpenStudy (anonymous):

the integral of (x/(4x^2+9))

OpenStudy (anonymous):

Let me give you a hint to start you off: U = 4x^2 + 9 dU = 8x dU/8 = x

OpenStudy (campbell_st):

this is a log function as the derivative of the denominator is 8x and the numerator is x... so you basically have \[\frac{f'(x)}{f(x)}\] so making an adjustment you have \[\frac{1}{8} \times \frac{8x}{4x^2 + 9}\] which is \[\frac{1}{8} \int\limits \frac{8x}{4x^2 + 9} dx = \frac{1}{8} \ln(4x^2 + 9) + C\] this can be simplified further by using log and index laws. hope this helps

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Medal? :D